VIBRATIONAL RELAXATION OF THE CH STRETCH FUNDAMENTAL IN CHBr_{3}

\underline{SAI G. RAMESH} and EDWIN L. SIBERT III, Department of Chemistry and Theoretical Chemistry Institute, University of Wisconsin-Madison, Madison, Wisconsin 53276.

In continuation of our work on haloforms, we have modelled the decay of CH stretch excitation in bromoform using molecular dynamics simulations. An intermolecular force field has been obtained by fitting ab initio energies at select CHBr_{3} dimer geometries to a potential function. The solvent forces on vibrational modes obtained in the simulation have been used to compute relaxation rates. The Landau-Teller approach points to a single acceptor state in the initial step of CH stretch relaxation. The timescale for this process has been calculated to be 50-90 ps, which agrees well with the experimental value of 50 ps. The reason for the selectivity of the acceptor will be elaborated. Results from a time dependent approach to the decay rates will also be discussed.